# DK-SLAM: Monocular Visual SLAM with Deep Keypoints Adaptive Learning, Tracking and Loop-Closing

### v1

@@ III-A4 Keypoint Distribution Strategy
Furthermore, we incorporate a keypoint distribution strategy during the feature extraction stage. We uniformly distribute the entire set of keypoints across a multi-scale image pyramid. Within each layer of the image pyramid, a predefined number of keypoints is distributed evenly within the grid of different regions on the image. This distribution strategy ensures a balanced dispersion of keypoints across all scales and corners within an image, preventing concentration in specific areas.
Figure 2: A diagram of our proposed Coarse-to-Fine Tracking.
III-B Coarse-to-Fine Keypoint Matching and Tracking
Keypoint matching accuracy is crucial for visual SLAM performance. Traditional methods like ORB-SLAM use a motion tracking module based on adjacent frames, assuming a uniform motion model. However, this assumption often fails, leading to difficulties in finding correct matching correspondences and complicating the optimization process.
@@ III-B1 Semi-direct Coarse Keypoint Tracking
Taking inspiration from the semi-direct method [3], we introduce a coarse tracking method based on photometric constraints. We assume that the light intensity and texture around matching keypoints are similar in adjacent frames. Map points from the last frame are projected onto the current frame using the relative pose. If the relative pose is accurate, the photometric loss of the projected keypoint’s patch will be minimal. As depicted in equations 5 and 6, the pixels of the keypoint’s patch, the map points in the last frame coordinate $\mathbf{p}^{l}_{i}$, and the relative pose of adjacent frames $\mathbf{\xi}^{\text{coarse}}_{lc}$ collectively form the photometric loss $L_{p}$. Given the known map points and patch pixel values of the last frame, we adjust the relative pose $\mathbf{\xi}^{\text{coarse}}_{lc}$ to minimize the photometric loss $L_{p}$.
@@ III-B2 Coarse-to-Fine Keypoint Tracking
The current frame’s pose serves as the vertex in the graph, with the number of vertices being subject to optimization. Map points from the current frame and the reprojection loss of these map points constitute the edges. We exclusively optimize the pose within the tracking module. As detailed in Equation 8, we iteratively refine the current frame pose $\hat{\mathbf{\xi}}^{c}_{w}$ to minimize the reprojection loss. The g2o [30] framework is employed for optimizing the pose graph.
@@
III-C Deep Keypoints based Loop Closing
Our DK-SLAM’s loop closure module includes detection and correction. To address poor generalization, we employ online learning-based Bag of Words (BoW) for the loop detector. Closed-loop nodes are identified, and the global map is optimized using the loop relative pose.
III-C1 Online learning based Binary Bow
Unlike handcrafted descriptors, deep descriptors occupy a larger feature space. Offline-trained BoW models may struggle to capture this space, making it difficult to distinguish deep feature indices in BoW leaf nodes. To tackle this, we introduce an online learning-based BoW that uses features exclusively from testing scene data. Constructing a tree-like structure with floating descriptors is challenging, so, inspired by [31], we employ binary hash transformation for processing deep features. As depicted in Equation 9, for descriptor vector values $\mathbf{d}_{i}$ less than 0, the processed $\mathbf{\hat{d}}_{i}$ is set to 0. For descriptor values greater than 0, the processed $\mathbf{\hat{d}}_{i}$ is set to 1, where $i$ ranges from 0 to 256.
Figure 3: Illustration of our proposed Online Learning based Binary BoW. The BoW is constructed incrementally, with matched descriptors in the keyframes database stored within the same leaf node. In the presence of unmatched descriptors in the current keyframe, a new leaf node is created.
III-C2 Loop Node Detection
As shown in the Figure 3, the loop closure thread receives keyframes filtered from the local mapping thread, storing them in the inverted database to construct a Bag-of-Words (BoW). The iBoW-LCD utilizes the feature descriptors of the current frame $K_{c}$ to match descriptors stored in the database. In iBoW-LCD, each keyframe is converted into a unique BoW vector for representation. We compute the similarity score between the current keyframe $K_{c}$ and the BoW vectors of inverted keyframes to identify the most similar candidate keyframe $K_{h}$. However, the candidate keyframe $K_{h}$ only considers differences in the numerical values of descriptors and neglects differences in descriptor positions, which may lead to mismatches.
@@ III-C2 Loop Node Detection
Initially, we employ the Grid-Based Motion Statistics (GMS) [32] based brute force matching method to match keypoints between candidate closed-loop nodes. If the number of matches is below 20, the candidate loop keyframe is classified as a mismatch. GMS operates on the principle that the number of keypoints near correctly matched feature points should surpass the number near incorrectly matched keypoints. Following a procedure similar to ORB-SLAM2, we compute the similarity transformation matrix $\mathbf{T}_{\text{sim}}$ between candidate closed-loop nodes through matched keypoints. Subsequently, $\mathbf{T}_{\text{sim}}$ is applied to match map points between the current keyframe $K_{c}$ and the candidate keyframe $K_{h}$. Further optimization of the similarity transformation matrix $\mathbf{T}_{sim}$ is conducted using the matched map points. If the number of inliers in the optimization function exceeds 20, the candidate keyframe $K_{h}$ is deemed a matched keyframe $K_{m}$.
III-C3 Global Map Correction by Loop Closure
@@ IV Experiments
TABLE I: The pose evaluation on the KITTI odometry dataset.
Method Sensor Metric 00 02 05 07 09 10 Avg
VISO-M Mono $t_{rel}$ 36.95 21.98 17.20 20.00 29.01 28.52 25.61
$r_{rel}$ 2.42 1.22 3.52 5.30 1.32 3.23 2.84
LDSO Mono $t_{rel}$ 2.84 4.91 2.07 12.76 23.14 9.20 9.15
$r_{rel}$ 0.36 0.74 0.24 5.13 0.21 0.21 1.15
ORB-SLAM3 Mono $t_{rel}$ 2.80 6.04 3.41 2.14 3.13 10.78 4.72
$r_{rel}$ 0.3 0.38 0.39 0.42 0.54 0.36 0.40
LIFT-SLAM Mono $t_{rel}$ 3.18 8.73 6.09 2.42 19.91 9.72 8.34
$r_{rel}$ 2.99 2.49 3.11 4.02 2.14 2.24 2.83
Ours Mono $t_{rel}$ 2.57 4.38 1.98 1.11 2.97 7.03 3.34
$r_{rel}$ 0.31 0.27 0.26 0.28 0.29 0.23 0.27
@@ IV-A2 KITTI Odometry Dataset
The KITTI odometry dataset [34] serves as a benchmark for self-driving scenarios, offering stereo RGB and grayscale images, along with Lidar and pose ground truth data. Grayscale images from sequences 00, 02, 05, 07, 09, and 10 are utilized for pose evaluation.
@@ IV-A3 EuRoc MAV Dataset
The EuRoc dataset [35] represents an indoor MAV flight dataset, encompassing stereo grayscale images, IMU data, and Leica Vicon’s pose ground truth. Grayscale images are collected at a frequency of 20Hz. Due to space limitations, we focus on sequences MH01-05 and V101 for SLAM performance evaluation.
IV-B Pose Evaluation on the KITTI Dataset
To assess our SLAM’s effectiveness, we experimented with the KITTI dataset using the official evaluation, computing root-mean-square error (RMSE) for translation and rotation vectors. The evaluation spanned sequences from 100m to 800m, providing an overall pose accuracy metric.
We thoroughly evaluated our SLAM’s tracking and loop closure in complex environments, focusing on KITTI’s 00, 02, 05, 07, 09, and 10 sequences. Pose estimation comparisons were made with LDSO [2], ORB-SLAM3 [12], VISO-M [36], and LIFT-SLAM [8]. VISO-M exhibited the lowest accuracy, facing failure scenarios, particularly in low-light conditions where corner detection proved unstable. LDSO, featuring a loop closure module, addressed accumulated errors in challenging scenes. ORB-SLAM3, an advanced keypoint-based SLAM with FAST corner detection and BRIFE descriptors, demonstrated exceptional performance across most scenes. Table I highlights ORB-SLAM3’s excellent performance in most sequences. However, a notable performance drop is observed in sequence 10, attributed to a lack of reliable corner points leading to unstable tracking and impacting overall pose estimation accuracy. Contrastingly, LIFT-SLAM, relying on learning-based local features, exhibits subpar performance across most sequences. This underscores that learning-based features may not universally surpass traditional features. ORB’s consideration of scale, direction, and a traditional SLAM technique for even keypoint distribution contribute to its competitive edge.
@@ IV-B Pose Evaluation on the KITTI Dataset
TABLE II: Comparison of absolute translation errors (in meters) between the proposed DK-SLAM and other baselines on the EuRoC dataset. Scaling with the ground truth is necessary for evaluation due to the absence of absolute scale in monocular visual SLAM methods, as noted in the table. ”-” indicates performance failure.
@@
Ours 0.013 0.013 0.027 0.077 0.055 0.0370
Learning-based local features perform well in matching benchmarks, but face challenges in SLAM due to correlated keypoints in dynamic scenes. Ignoring these correlations limits the exploitation of learning-based features’ advantages. Additionally, the feature space of learning-based features differs from handcrafted ones. Offline Bag-of-Words (BoWs) classify input features quickly, but traditional descriptors like BRIEF have a limited description space. Learning-based feature training is a ”black box,” complicating understanding, and testing performance depends on the training set’s distribution, hindering the acquisition of a generalized BoW.
Our proposed SLAM excels in most KITTI sequences, employing a coarse-to-fine matching strategy using keypoint-surrounding patches. Pose estimation optimization leverages the 3D-2D matching relationship. Online training constructs a learning-based feature BoW using prior experienced data, limiting the BoW’s feature description space. Results confirm the online BoW’s effectiveness in loop scene detection, with the loop closure module accurately rectifying cumulative errors.

### v2

@@ III-A4 The Distribution Strategy of Deep Keypoints
To ensure a balanced dispersion of learned keypoints across all scales and corners within an image, we incorporate a keypoint distribution strategy during the feature extraction stage. We uniformly distribute the entire set of keypoints across a multi-scale image pyramid. Within each layer of the image pyramid, a predefined number of keypoints is distributed evenly within the grid of different regions on the image. Thus this strategy can prevent the concentration of learned keypoints in specific areas.
Figure 2: Diagram of our proposed coarse-to-fine two-stage keypoint tracking strategy. This process begins with relative pose estimation through patch photometric loss optimization, followed by refinement using the 3D-2D keypoint relationship for enhanced accuracy.
III-B Coarse-to-Fine Keypoint Tracking
Accurate keypoint matching is essential for visual SLAM performance. Traditional systems like ORB-SLAM assume uniform motion between adjacent frames, often leading to challenges in finding correct matches and complicating optimization when this assumption fails. Unlike ORB features, which are straightforward low-level image descriptors, learned features generate complex high-dimensional embeddings, making matching and outlier filtering more difficult. To address these issues, we propose a two-stage deep keypoint tracking strategy to enhance matching robustness and accuracy.
@@ III-B1 Semi-direct Coarse Keypoint Tracking
Taking inspiration from the semi-direct visual odometry method [3], we introduce a coarse tracking method based on photometric constraints. We assume that the light intensity and texture around matching keypoints are similar in adjacent image frames. Map points from the last frame can be projected onto the current frame using the relative pose. If the relative pose is accurate, the photometric loss of the projected keypoint’s patch will be minimal. As depicted in Equations 5 and 6, the pixels of the keypoint’s patch, the image points in the last frame coordinate $\mathbf{p}^{l}_{i}$, and the relative pose of adjacent frames $\mathbf{\xi}^{\text{coarse}}_{lc}$ collectively form the photometric loss $L_{p}$. Given the known map points and patch pixel values of the last frame, we adjust the relative pose $\mathbf{\xi}^{\text{coarse}}_{lc}$ to minimize the photometric loss $L_{p}$.
@@ III-B2 Coarse-to-Fine Keypoint Tracking
The current frame’s pose serves as the vertex in the graph, with the number of vertices being subject to optimization. Map points from the current frame and the reprojection loss of these map points constitute the edges. We exclusively optimize the pose within the tracking module. As detailed in Equation 8, we iteratively refine the current frame pose $\hat{\mathbf{\xi}}^{c}_{w}$ to minimize the reprojection loss. In our implementation, the g2o [38] framework is employed for optimizing this pose graph.
@@
Figure 3: The illustration of our proposed Online Learning based Binary BoW. The BoW is constructed incrementally, with matched descriptors in the keyframes database stored within the same leaf node. In the presence of unmatched descriptors in the current keyframe, a new leaf node is created.
III-C Deep Keypoint Based Loop Closing
Our DK-SLAM system features a deep-keypoint-based loop closure module that performs both loop-closure detection and correction. To improve generalization in unfamiliar environments, we utilize an online learning-based Bag-of-Words (BoW) model for effective loop-closure detection. Once loop closure is detected, the system identifies the closed-loop nodes and optimizes the global map using the relative poses from the detected loops.
III-C1 Online Learning for Binary BoW
Unlike handcrafted descriptors, learned descriptors occupy a larger feature space, e.g. 128 dimensions in our case. Offline-trained BoW models may struggle to capture this space, making it difficult to distinguish deep feature indices in BoW leaf nodes. To tackle this, we introduce an online learning-based BoW model that uses extracted features exclusively from testing scene data. Constructing a tree-like structure with floating descriptors is challenging due to the large memory and time consumption, so, inspired by [39], we employ binary hash transformation for processing deep features. As depicted in Equation 9, for descriptor vector values $\mathbf{d}_{i}$ less than 0, the processed $\mathbf{\hat{d}}_{i}$ is set to 0. For descriptor values greater than 0, the processed $\mathbf{\hat{d}}_{i}$ is set to 1, where $i$ ranges from 0 to 256.
@@ III-C1 Online Learning for Binary BoW
Figure 5: Mapping results generated by our proposed DK-SLAM system.
TABLE I: The pose evaluation on the KITTI dataset: our DK-SLAM system outperforms both representative traditional and learning-based monocular SLAM methods. Sequence 01 is excluded from the evaluation due to its inherent challenges for all monocular SLAM baselines, which prevent the generation of reasonable comparative results.
Method Sensor Metric 00 02 03 04 05 06 07 08 09 10 Avg
VISO-M Mono $t_{rel}$ 36.95 21.98 16.14 2.61 17.20 7.91 20.00 39.78 29.01 28.52 22.01
$r_{rel}$ 2.42 1.22 2.67 1.53 3.52 1.83 5.30 1.99 1.32 3.23 2.50
LDSO Mono $t_{rel}$ 2.84 4.91 2.75 1.58 2.07 6.49 12.76 30.17 23.14 9.20 9.59
$r_{rel}$ 0.36 0.74 0.16 0.19 0.24 0.16 5.13 0.24 0.21 0.21 0.76
ORB-SLAM3 Mono $t_{rel}$ 2.80 6.04 2.71 2.22 3.41 7.07 2.14 10.35 3.13 10.78 5.07
$r_{rel}$ 0.3 0.38 0.17 0.24 0.39 0.18 0.42 0.31 0.54 0.36 0.33
LIFT-SLAM Mono $t_{rel}$ 3.18 8.73 1.46 2.22 6.09 12.24 2.42 47.10 19.91 9.72 11.31
$r_{rel}$ 2.99 2.49 0.34 0.48 3.11 2.91 4.02 2.02 2.14 2.24 2.27
DK-SLAM (Ours) Mono $t_{rel}$ 2.57 4.38 4.70 1.12 1.98 8.20 1.11 7.61 2.97 7.03 4.17
$r_{rel}$ 0.31 0.27 0.17 0.24 0.26 0.17 0.28 0.28 0.29 0.23 0.25
III-C2 Loop Node Detection
As shown in the Figure 3, the loop closure thread receives keyframes filtered from the local mapping thread, storing them in the inverted database to construct a Bag-of-Words (BoW) model. The BoW model utilizes the feature descriptors of the current frame $K_{c}$ to match descriptors stored in the database. Each keyframe is converted into a unique BoW vector for representation. We compute the similarity score between the current keyframe $K_{c}$ and the BoW vectors of inverted keyframes to identify the most similar candidate keyframe $K_{h}$. However, the candidate keyframe $K_{h}$ only considers differences in the numerical values of descriptors and neglects differences in descriptor positions, which may lead to mismatches.
Thus, we first employ the Grid-Based Motion Statistics (GMS) [40] based brute force matching method to match deep keypoints between candidate closed-loop nodes. If the number of matches is below a threshold (i.e. 20 in our system), the candidate loop keyframe is classified as a mismatch. GMS operates on the principle that the number of keypoints near correctly matched feature points should surpass the number near incorrectly matched keypoints. Following a procedure similar to ORB-SLAM2[4], we compute the similarity transformation matrix $\mathbf{T}_{\text{sim}}$ between candidate closed-loop nodes through matched keypoints. Subsequently, $\mathbf{T}_{\text{sim}}$ is applied to match map points between the current keyframe $K_{c}$ and the candidate keyframe $K_{h}$. Further optimization of the similarity transformation matrix $\mathbf{T}_{sim}$ is conducted using the matched map points. If the number of inliers in the optimization function exceeds the threshold, the candidate keyframe $K_{h}$ is deemed a matched keyframe $K_{m}$.
III-C3 Global Map Correction via Loop-Closing
@@ IV Experiments
This section discusses the implementation details and evaluates the pose estimation performance of our proposed DK-SLAM on two widely adopted public datasets: the KITTI dataset, representing the car-driving scenario, and the EuRoC dataset, representing drone navigation. Moreover, we conducted an extensive ablation study to validate the effectiveness of our proposed key modules, including learned feature extraction, coarse-to-fine matching strategy, and online learning-based Bag-of-Words (BoW).
@@ IV-A2 KITTI Odometry Dataset
The KITTI odometry dataset [42] serves as a benchmark for self-driving scenarios, offering stereo RGB and grayscale images, along with Lidar and pose ground truth data. Grayscale images from sequences 00, 02, 03, 04, 05, 06, 07, 08, 09, and 10 are utilized for pose evaluation. Sequence 01 is excluded from the evaluation due to its inherent challenges for monocular SLAM systems.
@@ IV-A3 EuRoC MAV Dataset
The EuRoC dataset [43] represents an indoor MAV flight dataset, encompassing stereo grayscale images, IMU data, and Leica Vicon’s pose ground truth. Grayscale images are collected at a frequency of 20Hz. Due to space limitations, we focus on sequences MH01-05 for SLAM performance evaluation.
IV-B Pose Evaluation in the Car-Driving Scenario
To evaluate the effectiveness of our SLAM system in the car-driving scenario, we conducted experiments using the KITTI dataset and employed the official evaluation metrics, computing root-mean-square error (RMSE) for both translation and rotation vectors. This evaluation spanned sequences ranging from 100m to 800m, providing an overall metric for pose accuracy. We conducted a comprehensive assessment of our SLAM system’s tracking and loop closure capabilities in complex environments, focusing on Sequences 00, 02, 03, 04, 05, 06, 07, 08, 09, and 10 from the KITTI dataset. We excluded Sequence 01 from our evaluation due to its unique challenges for monocular SLAM systems, as a highway scene with few distinctive feature points, especially near the horizon, it led to significant scale degradation in monocular SLAM performance.
We benchmarked our DK-SLAM against several leading SLAM systems: LDSO [44], ORB-SLAM3 [5], VISO-M [45], and LIFT-SLAM [9]. Among these, VISO-M exhibited the least accuracy, particularly struggling in low-light conditions where its corner detection algorithms were unstable. LDSO, which includes a loop closure module, managed to reduce accumulated errors effectively in challenging scenarios. ORB-SLAM3, a sophisticated keypoint-based SLAM system that uses FAST corner detection and BRIEF descriptors, performed exceptionally well across most sequences as shown in Table I. However, it showed a noticeable drop in performance on Sequence 10, primarily due to a scarcity of reliable corner points, which resulted in unstable tracking and adversely affected overall pose estimation accuracy. The reliance on BRIEF, which describes local photometric information around corners, made ORB-SLAM3 particularly sensitive to lighting changes, thereby impacting feature tracking performance. In environments with significant lighting variations, the inability to track reliable feature points at the front-end increased positioning errors. Although the back-end loop closure could correct some accumulated errors, it struggled to improve overall accuracy. In contrast, LIFT-SLAM, which depends on learning-based local features, generally underperformed across most sequences. This outcome suggests that while learning-based features have potential, they do not consistently outperform traditional feature-based methods. ORB-SLAM3’s attention to scale, direction, and a traditional SLAM system for achieving even keypoint distribution gave it a competitive advantage, helping to filter out mismatched features at the front-end.
Our DK-SLAM system demonstrated exceptional performance across most KITTI sequences. Specifically, DK-SLAM shows a substantial improvement over traditional monocular ORB-SLAM3, achieving approximately 17.7% better translation accuracy and 24.2% higher rotation accuracy on the KITTI dataset. When compared to LIFT-SLAM, a leading SLAM system based on learned features, DK-SLAM surpasses it by nearly 2.7 times in translation accuracy and 9 times in rotation accuracy. This significant enhancement in performance is attributed to DK-SLAM’s use of meta learning based feature extraction and a coarse-to-fine matching strategy, which focuses on keypoint-surrounding patches and optimizes pose estimation through precise 3D-2D matching relationships. Additionally, our system’s online training capability builds a learning-based Bag-of-Words (BoW) model using previously acquired data, efficiently constraining the BoW’s feature description space and improving loop scene detection. This approach enables the loop closure module to accurately correct cumulative errors. Moreover, as illustrated in Figure 5, DK-SLAM excels in mapping performance, providing detailed geometric insights. The incorporation of MAML-based feature training enhances the generalization capability of the feature extractor, allowing it to capture more comprehensive scene details. This combination of strategies ensures that DK-SLAM offers robust and accurate performance in varied and complex environments.
(a) Sequence MH01
(b) Sequence MH02
(c) Sequence MH03
(d) Sequence MH04
(e) Sequence MH05
Figure 6: The generated trajectories of our proposed DK-SLAM on the Sequence MH01, MH02, MH03, MH04 and MH05 of the EuRoC dataset, comparing with DSO and ORB-SLAM3.
@@ IV-B Pose Evaluation in the Car-Driving Scenario
TABLE II: Comparison of absolute translation errors (in meters) between the proposed DK-SLAM and other baselines on the EuRoC dataset. Scaling with the ground truth is necessary for evaluation due to the absence of absolute scale in monocular visual SLAM methods, as noted in the table. "-" indicates performance failure.
@@
DK-SLAM 0.013 0.013 0.027 0.077 0.055 0.0370
TABLE III: The ablation study into the key modules in our DK-SLAM system.
Method Meta Learning Two-Stage Tracking Metric 00 02 03 04 05 06 07 08 09 10 Avg
Ours1 $\checkmark$ $t_{rel}$ 2.88 - 4.29 1.24 2.80 7.64 1.76 - 2.70 - -
$r_{rel}$ 0.41 - 0.16 0.31 0.35 0.18 0.64 - 0.24 - -
Ours2 $\checkmark$ $t_{rel}$ 2.92 4.42 4.54 2.01 2.48 7.97 2.11 8.85 4.03 8.48 4.78
$r_{rel}$ 0.38 0.34 0.21 0.16 0.34 0.18 0.46 0.26 0.20 0.23 0.28
Ours3 $\checkmark$ $\checkmark$ $t_{rel}$ 2.57 4.38 4.70 1.12 1.98 8.20 1.11 7.61 2.97 7.03 4.17
$r_{rel}$ 0.31 0.27 0.17 0.24 0.26 0.17 0.28 0.28 0.29 0.23 0.25
IV-C Pose Evaluation in the UAV Scenario
As shown in Table II and Figure 6, we come to validate our proposed DK-SLAM on the EuRoC dataset. Following the previous research, we address the scale ambiguity with umeyama alignment [46] for monocular vision methods. Comparisons include ORB-SLAM3 [5], DSO [2], DSM [47], SVO [3], and LIFE-SLAM [9]. Results, derived from [7] for ORB-SLAM3, DSO, DSM, and SVO, and [9] for LIFE-SLAM. We utilize the absolute translation error as the evaluation metric.
The evaluation results reveal that the learned feature based SLAM, i.e. LIFE-SLAM, often fails in drone scenarios, as it relies on LIFT as a binary feature extractor and uses the Hamming distance between keypoints in consecutive frames for matching. In addition, LIFE-SLAM does not consider the structural relationships between keypoints, leading to a significant number of mismatches and subsequent feature tracking failures. ORB-SLAM3 [5] employs an offline-trained Bag-of-Words (DBow3) for loop closure. Due to the sensitivity of its handcrafted descriptors to lighting variations, ORB-SLAM3 struggles with lower loop detection accuracy in complex lighting conditions, which diminishes the overall performance of the system.
In contrast, our DK-SLAM demonstrates outstanding localization performance. It uses a two-stage tracking strategy that precisely locates feature points and minimizes incorrect matches. The learned features in DK-SLAM are robust, performing well even in low-light conditions and supporting stable feature tracking. Moreover, the online learning-based deep Bag-of-Words model excels in loop detection. This is particularly evident in the MH02 sequence, where DK-SLAM surpasses ORB-SLAM3. This superior performance is due to the use of learned features for constructing the Bag-of-Words model, which provides detailed scene differentiation, and the online training capability, which allows for rapid adaptation to changing environments.